A purse containing only P5 and P10 coins, has 6 more P10 coins than P5 coins. The total value of the money is P360, how many P5 coins and P10 coins are in the purse.
P.S.
i need the formulation also. tnx
: A purse containing only P5 and P10 coins, has 6 more P10 coins than P5 coins. The total value of the money is P360, how many P5 coins and P10 coins are in the purse.
P.S.
i need the formulation also. tnx

Answer by jojo14344(888)


Total number of coins P5=x
Total number of coins P10= x+6, six more than P5 coins
Remember:
When total number of P5 coins multiplied by P5 + total number of P10 coins multplied by P10 equals P360.Putting into eqn:
P5x+P10(x+6)=P360, WORKING EQN
P5+P10x+P60=P360, distribute & combine similar terms then add/subtract
P15x=P360-P60=P300
cross(P15)x/cross(P15)=cross(P300)20/cross(P15), divide P15 to isolate x
x=20 -------------------------> number of P5 coins
20+6=26 ----------------------> number of P10 coins
Check via WORKING EQN,
P5*20+P10(20+6)=P360
P100+P260=P360
P360=P360
Thank you,
Jojo